OpenStudy (anonymous):

How do I use the square root on my calculator?? I am trying to solve 3v- 5v+2=0 ?? The 3v is squared.

OpenStudy (amistre64):

3v * 3v

OpenStudy (amistre64):

square root of 3v? or 3v squared?

OpenStudy (amistre64):

i get it, your trying to use the quadratic formula

OpenStudy (amistre64):

\[\frac{5\pm\sqrt{25-4(2)(3)}}{2(3)}\] \[\frac{5\pm\sqrt{25-24}}{2(3)}\] \[\frac{5\pm\sqrt{1}}{6}\] \[\frac{5\pm1}{6}\]

OpenStudy (amistre64):

not everything needs to be done by a calculator .....

OpenStudy (anonymous):

3v*-5v+2=0 I am to solve by factoring.. I get stuck at getting the square root.

OpenStudy (amistre64):

you dont need to get a square root to solve by factoring that im aware of

OpenStudy (anonymous):

it hink there is no need for using quadratic formula. . the given quadratic equation is factorable. . here, , (3v-2)(v-1) . therefor, zeroes are 2 and 1

OpenStudy (anonymous):

I am aware I can start it.. I just can't finish it off

OpenStudy (amistre64):

2 is a zero?

OpenStudy (anonymous):

???

OpenStudy (amistre64):

jerwyn made a typo; 2/3 is a zero not 2 :)

OpenStudy (anonymous):

ahh, yeah. . @amistre

OpenStudy (anonymous):

How can I perform the last step using my calculator/?? that is what I need to show..